DsneadBeats:

In ΔQRS, q = 1.2 cm, r = 9.7 cm and s=9.2 cm. Find the measure of ∠Q to the nearest degree.

surjithayer:

\[\cos Q=\frac{ 9.7^2+9.2^2-1.2^2 }{ 2\times 9.7 \times 9.2 }\] \[\cos Q=\frac{ 94.09+84.64-1.44 }{ 178.48}\] \[\cos Q=\frac{ 177.24 }{ 178.48 }\] \[Q=\cos ^{-1}(\frac{ 177.24 }{ 178.48})\approx 6.76^\circ \approx 7 ^\circ \]

surjithayer:



curriful:

the drawing is insane